# Restoring the Forecasting Power of Google Trends with Statistical Preprocessing

### v1

@@ 3.1 Missing Values and Sampling Variability
Hierarchical clustering (Ward 1963) is a data analysis technique that creates a sequence of nested clusters based on a hierarchical structure (Maharaj et al. 2019). We use Ward’s method to determine how clusters are merged, as it minimizes the total within-cluster variance at each step. This approach produces spherical clusters of relatively similar sizes (Maharaj et al. 2019, Everitt et al. 2011), ensures clear separation, and is regarded as one of the most robust and effective methods for noisy data (Balcan et al. 2014). Clusters with high similarity are merged together in successive steps. In hierarchical clustering, smaller, closely related clusters can be combined into broader groups, while larger ones can be subdivided to enhance relevance and specificity.
[…]
To determine the optimal number of clusters for hierarchical clustering, we use the elbow method, which plots the within-cluster sum of squares (WCSS) against the number of clusters. WCSS, defined as $\sum_{i=1}^{k}\sum_{x\in C_{i}}|x-\mu_{i}|^{2}$, where $\mu_{i}$ is the average time series of cluster $C_{i}$, measures the sum of squared distances between each time series and its cluster center (Maharaj et al. 2019, Hastie et al. 2017). As the number of clusters increases, WCSS decreases, with the optimal point occurring at the “elbow,” where adding more clusters yields diminishing returns. At the elbow, clusters strike a balance between capturing meaningful patterns and maintaining a sufficient size to surpass Google’s privacy thresholds. If a single cluster remains disproportionately large compared to the others, we apply a second round of hierarchical clustering to further subdivide it. The categorized queries are then grouped using the Boolean “+” operator, and their combined search volumes are retrieved from Google Trends. We finalize the clusters after confirming they contain low percentages of zeros.
Clustering not only addresses missing values, but also sampling variability. Popular keywords with high search volumes exhibit consistent behavior across different samples, while less popular terms with sparse data are subject to higher sampling variability (Medeiros & Pires 2021). Grouping such queries into clusters increases search volumes, effectively reducing hypergeometric sampling randomness and improving data reliability.
@@ 3.2 Noise
Noise in Google Trends data arises from random sampling variations, which can obscure the underlying signal and lead to overfitting. While various statistical techniques can estimate the true signal, real-time forecasting requires smoothing time series using only past data to prevent information leakage, where data not available at the time of estimation is inadvertently included in the model (Joseph 2022, Borup & Schütte 2022). Exponential smoothing and moving averages rely on past values but may introduce phase shifting, while K-nearest neighbors, kernel smoothing, and LOESS require both past and future points. Polynomial regression risks overfitting or underfitting, an Fourier transforms and wavelets need the entire time series for frequency and time-frequency analysis (Shumway et al. 2006, Brockwell & Davis 2016). Such constraints make these methods unsuitable for real-time applications. To overcome these challenges, we apply smoothing splines iteratively within a rolling window of past values. This approach offers localized data fitting, relies exclusively on past observations, provides precise control over smoothing, and has been shown to effectively remove noise (Musial et al. 2011, Navarrete & Viswanath 2018).
Smoothing splines (Reinsch 1967) aim to minimize the function
[…]
where $x_{t}$ are the observations, $f_{t}$ is the cubic spline, and $\lambda$ is the smoothing parameter controlling the degree of smoothness (Hastie et al. 2017, Shumway et al. 2006). Larger $\lambda$ values produce smoother time series but risk shrinking peaks and flattening the signal, while smaller values may retain excessive noise. To balance these effects and avoid over-smoothing, we restrict $\lambda$ to the range [0.1, 2].
We implement an iterative denoising process tailored for real-time settings, performing one-step-ahead predictions of Google Trends search volumes. The data is split into training and testing sets, with optimal $\lambda$ values for each keyword identified through a grid search on the training set, and the test set used for out-of-sample denoising. During training, we fit smoothing splines on a rolling window of recent observations, and the next data point outside the window is predicted. The window then advances by one observation, and the process is repeated across the training set. During testing, smoothing splines are applied similarly, but the last fitted observation within the rolling window, not the predicted one, is used as the denoised value. This approach prevents overfitting to recent trends and ensures smoother transitions. The window size is determined through visualizations and performance evaluations.
@@
Denoising is applied only to time series with high noise levels, identified by evaluating the RMSE associated with the optimal $\lambda$. Higher RMSE values indicate that noise dominates the series, making it difficult for the model to produce accurate predictions. If the RMSE for a series exceeds a predefined threshold, determined through analysis of time series patterns, smoothing is necessary. This strategy allows us to selectively improve noisy series while avoiding unnecessary adjustments for cleaner ones.
@@ 3.3 Trend
Trends in Google Trends data arise from long-term changes in search volumes, often driven by shifts in public interest or algorithm updates. While most keywords have search volumes with a relatively stable mean over time, some exhibit persistent trends. However, for applications such as predicting infectious disease surges, unemployment rates, or stock prices, the focus is typically on short-term patterns. Long-term trends can obscure these dynamics and introduce noise into forecasting models. Furthermore, a time series with a trend is non-stationary, complicating the accurate estimation of autocorrelations between observations (Shumway et al. 2006) and potentially distorting variable relationships. We apply methods such as model fitting or differencing, to remove trends and restore stationarity. Having access to Google Trends data since 2004 provides an advantage, allowing for a clear estimation of the long-term trend.
Time series trends can be classified as either deterministic or stochastic. A deterministic trend is a fixed and predictable pattern over time, following a linear or quadratic shape, and can be removed through detrending to achieve stationarity. A stochastic trend, on the other hand, arises from the cumulative effect of past values and random shocks, and should be eliminated through differencing. To identify the presence and the type of trend in a time series, we apply the Augmented Dickey-Fuller (ADF) test (Dickey & Fuller 1979) to each search volume time series. The null hypothesis ($H_{0}$) states that the time series has a unit root and is non-stationary, while the alternative hypothesis ($H_{1}$) is that the time series is stationary, potentially around a deterministic trend (Joseph 2022). Since search volumes are strictly non-negative, their time series have a non-zero mean. The ADF test fits a regression model to the time series data, as follows (Valenzuela et al. 2019):
@@ 4.1 Evaluation of Clustering
Figure 5 analyzes the percentage of missingness before and after clustering for all 28,580 keywords in our dataset. 814 queries with $\leq 30\%$ of zeros were retained in their original form, 8,841 with between 30% and 99% zeros were grouped into 341 clusters, and the remainder were discarded. Figure 5(a) shows that these individual search volumes often contain sparse data with many zeros. After clustering, the percentage of zeros in the aggregated search volumes is notably reduced, as shown in Figure 5(b). Most clusters fall within the [0–10]range of zeros, with very few clusters in the [90–99]range, demonstrating a substantial reduction in sparsity. Our method thus transforms sparse data dominated by zeros into denser aggregated data, improving signal quality.
@@ 4.2 Evaluation of Denoising
Effective denoising should minimize short-term fluctuations while preserving underlying patterns. The Signal-to-Noise Ratio (SNR) measures the power of the signal to the power of the noise (Palani & Kalaiyarasi 2022). Since real-world data lacks a clear distinction between signal and noise, Smith et al. (1999) define SNR as the ratio of the mean to the standard deviation of the data, $\text{SNR}=\mu/\sigma$, where $\mu$ reflects the overall magnitude of the signal, and $\sigma$ approximates the noise level. A higher SNR indicates a cleaner signal. We estimate the SNR at each time step as $\hat{\mu}_{t}/\hat{\sigma}_{t}$, where $\hat{\mu}_{t}$ and $\hat{\sigma}_{t}$ are the mean and standard deviation of the time series at time $t$ across downloads. The log ratio of SNR between denoised and raw data is then computed as:
log Ratio of $\text{log Ratio of SNR}=\log\left(\frac{\hat{\mu}_{t}^{\text{(denoised)}}}{\hat{\sigma}_{t}^{\text{(denoised)}}}\Big{/}\frac{\hat{\mu}_{t}^{\text{(raw)}}}{\hat{\sigma}_{t}^{\text{(raw)}}}\right).$ (3)
Figure 8 presents the log ratio of SNR across twenty-seven downloads. Figure 8(a) illustrates its temporal evolution for a single time series, while Figure 8(b) displays the results across all keywords. Predominantly positive values indicate that denoised data have higher SNR than raw data, confirming that denoising enhances signal quality by reducing noise and sampling variability.
@@ 4.3 Evaluation of Detrending
Figure 10 summarizes the distribution of $R^{2}$ across all keywords. Fitted trends in the raw data exhibit high $R^{2}$ values (mean $=0.77$), confirming that they were well estimated. After detrending, $R^{2}$ values decrease significantly (mean $=0.11$), demonstrating the effective removal of deterministic trends. There are a few outliers with low $R^{2}$ values before detrending and high values afterward, suggesting occasional misestimation, as stationarity is detected based on the training set and may not fully capture time series patterns.
@@ 5 Application to Influenza Prediction
To validate the effectiveness of our methodology, we use a univariate ARIMAX(1,1,1) model to compare the forecasting performance of our preprocessed Google Trends data with that of raw data in the context of real-time flu forecasting. ARIMA models are widely used for short-term predictions due to their ease of implementation, computational efficiency, and interpretability. In influenza forecasting, ARIMA has consistently demonstrated strong accuracy for 1-week-ahead predictions (Meyer et al. 2025, Mathis et al. 2024), making it a reliable baseline for evaluating more complex methods (Kandula & Shaman 2019). Studies have also shown that incorporating exogenous variables, such as Google Trends data, into autoregressive models significantly improves nowcasting performance (Lampos et al. 2015). Accordingly, we extend ARIMA with Google Trends data to evaluate the impact of our preprocessing approach on forecasting flu hospitalizations.
5.1 ARIMAX as a Predictive Model
AutoRegressive Integrated Moving Average with eXogenous variables (ARIMAX) (Tiao & Box 1981) is an extension of the ARIMA model (Box & Jenkins 1970) that incorporates external predictors. Let $y_{t}$ be the target variable and $x_{t,k}$ the $k^{\text{th}}$ exogenous variable at time $t$. The model used in this study is a univariate ARIMAX(1,1,1) with first-order differencing, expressed as:
$\Delta y_{t}=\mu+\phi_{1}\Delta y_{t-1}+\theta_{1}\epsilon_{t-1}+\epsilon_{t}+\sum_{k=1}^{K}\beta_{k}x_{t,k}.$ (4)
Here, $\Delta y_{t}$ denotes the differenced target series, $\mu$ is the mean of the differenced series, $\phi_{1}$ and $\theta_{1}$ represent the coefficients for the autoregressive (AR) and moving average (MA) terms, respectively, and $\epsilon_{t}$ is the current error term. The coefficients $\beta_{k}$ correspond to the exogenous variables (Peixeiro 2022). Readers unfamiliar with the general ARIMA model are referred to Supporting Information Section § B.
Model (4) is applied to each location for retrospective out-of-sample 1-week-ahead forecasts to simulate real-time predictions. This approach allows for comparing forecasting performance across three scenarios: using preprocessed Google Trends data, non-preprocessed data, and no exogenous variables. The model is trained on a rolling window of 104 weeks, with optimal parameters determined by minimizing the training loss, calculated as the sum of squared fitting errors within each training window. After each forecast, the rolling window shifts forward by one week, and the model is retrained to predict the following week’s target. To assess forecasting performance, we calculate the Mean Squared Error (MSE) between the ground truth and predicted values, $\text{MSE}=\frac{1}{T}\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}$, over the test period $t=1,...,T$. We then compute the Relative Efficiency (RE) as the ratio of MSEs between two models. The overall RE is based on the average MSE across all $n$ states as $\frac{\sum_{i=1}^{n}\text{MSE}_{i}\text{(ARIMAX(1,1,1))}}{\sum_{i=1}^{n}\text{MSE}_{i}\text{(ARIMA(1,1,1))}}$ (Ning et al. 2019), where ARIMA(1,1,1) without exogenous variables serves as the benchmark.

### v2

@@ 4.1 Missing Values and Sampling Variability
Hierarchical clustering (Ward, 1963) is a data analysis technique that creates a sequence of nested clusters based on a hierarchical structure (Maharaj et al., 2019). We use Ward’s method to determine how clusters are merged, as it minimizes the total within-cluster variance at each step. This approach produces spherical clusters of relatively similar sizes (Maharaj et al., 2019; Everitt et al., 2011), ensures clear separation, and is regarded as one of the most robust and effective methods for noisy data (Balcan et al., 2014). Clusters with high similarity are merged together in successive steps. In hierarchical clustering, smaller, closely related clusters can be combined into broader groups, while larger ones can be subdivided to enhance relevance and specificity.
The similarity metric used in our hierarchical clustering is correlation distance, which is defined as (1correlation matrix)(1-$\text${correlation matrix}). Since different phrasings of a query often produce similar search results, their search volumes tend to align as well. Correlation matches time series that share similar shapes and have synchronized peaks and troughs. It also enables sparse signals with relevant spikes to be grouped with more complete ones, provided their key patterns align. Clustering keywords based on correlations thus forms groups that not only share semantic meanings but also have combined search volumes that surpass Google’s privacy thresholds. This approach effectively addresses the challenge of zeros while preserving meaningful search activity.
To determine the optimal number of clusters for hierarchical clustering, we use the elbow method, which plots the within-cluster sum of squares (WCSS) against the number of clusters. WCSS, defined as $\sum_{i=1}^{k}\sum_{x\in C_{i}}|x-\mu_{i}|^{2}$, where $\mu_{i}$ is the average time series of cluster $C_{i}$, measures the sum of squared distances between each time series and its cluster center (Maharaj et al., 2019; Hastie et al., 2017). As the number of clusters increases, WCSS decreases, with the optimal point occurring at the “elbow,” where adding more clusters yields diminishing returns. At the elbow, clusters strike a balance between capturing meaningful patterns and maintaining a sufficient size to surpass Google’s privacy thresholds. If a single cluster remains disproportionately large compared to the others, we apply a second round of hierarchical clustering to further subdivide it. The categorized queries are then grouped using the Boolean “+” operator, and their combined search volumes are retrieved from Google Trends. We finalize the clusters after confirming they contain low percentages of zeros.
Clustering not only addresses missing values, but also sampling variability. Popular keywords with high search volumes exhibit consistent behavior across different samples, while less popular terms with sparse data are subject to higher sampling variability (Medeiros and Pires, 2021). Grouping such queries into clusters increases search volumes, effectively reducing hypergeometric sampling randomness and improving data reliability.
@@ 4.2 Noise
Noise in Google Trends data arises from random sampling variations, which can obscure the underlying signal and lead to overfitting. While various statistical techniques can estimate the true signal, real-time forecasting requires smoothing time series using only past data to prevent information leakage, where data not available at the time of estimation is inadvertently included in the model (Joseph, 2022; Borup and Schütte, 2022). Exponential smoothing and moving averages rely on past values but may introduce phase shifting, while K-nearest neighbors, kernel smoothing, and LOESS require both past and future points. Polynomial regression risks overfitting or underfitting, and Fourier transforms and wavelets need the entire time series for frequency and time-frequency analysis (Shumway et al., 2006; Brockwell and Davis, 2016). Such constraints make these methods unsuitable for real-time applications. To overcome these challenges, we apply smoothing splines iteratively within a rolling window of past values. This approach offers localized data fitting, relies exclusively on past observations, provides precise control over smoothing, and has been shown to effectively remove noise (Musial et al., 2011; Navarrete and Viswanath, 2018).
Smoothing splines (Reinsch, 1967) aim to minimize the function
$\sum_{t=1}^{n}(x_{t}-f_{t})^{2}+\lambda\int(f^{\prime\prime}_{t})^{2}dt,$ (1)
where $x_{t}$ are the observations, $f_{t}$ is the cubic spline, and $\lambda$ is the smoothing parameter controlling the degree of smoothness (Hastie et al., 2017; Shumway et al., 2006). Larger $\lambda$ values produce smoother time series but risk shrinking peaks and flattening the signal, while smaller values may retain excessive noise. To balance these effects and avoid over-smoothing, we restrict $\lambda$ to the range [0.1, 2].
We implement an iterative denoising process tailored for real-time settings, performing one-step-ahead predictions of Google Trends search volumes. The data is split into training and testing sets, with optimal $\lambda$ values for each keyword identified through a grid search on the training set, and the test set used for out-of-sample denoising. During training, we fit smoothing splines on a rolling window of recent observations, and the next data point outside the window is predicted. The window then advances by one observation, and the process is repeated across the training set. During testing, smoothing splines are applied similarly, but the last fitted observation within the rolling window, not the predicted one, is used as the denoised value. This approach prevents overfitting to recent trends and ensures smoother transitions.
@@
The rolling window length is chosen based on the characteristics of the data and domain knowledge. According to the CDC, an influenza season typically spans from October to May, or roughly 30 weeks (Centers for Disease Control and Prevention, 2025a, 2024). To avoid smoothing across multiple seasons and diluting within-season dynamics, the window should not exceed this activity period. We use a 20-week window, which represents about 70% of the season length, to ensure that both seasonal patterns and sudden spikes are retained. If the window is too short, denoising may not effectively reduce noise, while if it is too long, denoising may attenuate meaningful short-term variations.
Denoising is applied only to time series with high noise levels, identified by evaluating the training RMSE associated with the optimal $\lambda$. Higher RMSE values indicate that noise dominates the series, making it difficult for the model to produce accurate predictions. To determine which series to denoise, we set a threshold based on the empirical distribution of RMSEs across all time series. Specifically, series with RMSE greater than the median RMSE are considered noisy and are smoothed. This strategy allows us to selectively improve noisy series while avoiding unnecessary adjustments for cleaner ones.
@@ 4.3 Trend
Trends in Google Trends data arise from long-term changes in search volumes, often driven by shifts in public interest, dynamic normalization, or algorithm updates. While most keywords have search volumes with a relatively stable mean over time, some exhibit persistent trends. However, for applications such as predicting infectious disease surges, unemployment rates, or stock prices, the focus is typically on short-term patterns. Long-term trends can obscure these dynamics and introduce noise into forecasting models. Furthermore, a time series with a trend is non-stationary, complicating the accurate estimation of autocorrelations between observations (Shumway et al., 2006) and potentially distorting variable relationships. Detrending therefore enhances the interpretability of the data and supports more robust comparisons between search interest and target outcomes. We apply methods such as model fitting or differencing, to remove trends and restore stationarity. Having access to Google Trends data since 2004 provides an advantage, allowing for a clear estimation of the long-term trend.
Time series trends can be classified as either deterministic or stochastic. A deterministic trend is a fixed and predictable pattern over time, following a linear or quadratic shape, and can be removed through detrending to achieve stationarity. A stochastic trend, on the other hand, arises from the cumulative effect of past values and random shocks, and should be eliminated through differencing. To identify the presence and the type of trend in a time series, we apply the Augmented Dickey-Fuller (ADF) test (Dickey and Fuller, 1979) to each search volume time series. The null hypothesis ($H_{0}$) states that the time series has a unit root and is non-stationary, while the alternative hypothesis ($H_{1}$) is that the time series is stationary, potentially around a deterministic trend (Joseph, 2022). Since search volumes are strictly non-negative, their time series have a non-zero mean. The ADF test fits a regression model to the time series data, as follows (Valenzuela et al., 2019):
@@ 5.1 Evaluation of Clustering
Figure 6 analyzes the percentage of missingness before and after clustering for all 28,580 keywords in our dataset. 814 queries with $\leq 30\%$ of zeros were retained in their original form, 8,841 with between 30% and 99% zeros were grouped into 341 clusters, and the remainder were discarded. Figure 6(a) shows that these individual search volumes often contain sparse data with many zeros. After clustering, the percentage of zeros in the aggregated search volumes is notably reduced, as shown in Figure 6(b). Most clusters fall within the [0–10]% range of zeros, with very few clusters in the [90–99]% range, demonstrating a substantial reduction in sparsity. Our method thus transforms sparse data dominated by zeros into denser aggregated data, improving signal quality.
@@ 5.2 Evaluation of Denoising
Effective denoising should minimize short-term fluctuations while preserving underlying patterns. The Signal-to-Noise Ratio (SNR) measures the power of the signal to the power of the noise (Palani and Kalaiyarasi, 2022). Since real-world data lacks a clear distinction between signal and noise, Smith et al. (1999) define SNR as the ratio of the mean to the standard deviation of the data, $\text{SNR}=\mu/\sigma$, where $\mu$ reflects the overall magnitude of the signal, and $\sigma$ approximates the noise level. A higher SNR indicates a cleaner signal.
We downloaded repeated instances of the same query at each time point $t$, where each instance is an independent and identically distributed (iid) realization of the underlying signal plus Google-injected noise. Averaging these instances, referred to as signal averaging, reduces the effect of random fluctuations, as noise tends to cancel out when averaged, while the underlying signal is preserved (Dempster, 2001).
We estimate the SNR at each time step as $\text{SNR}_{t}=\frac{\text{mean(repeated)}}{\text{sd(repeated)}}=\hat{\mu}_{t}/\hat{\sigma}_{t}$, where $\hat{\mu}_{t}$ and $\hat{\sigma}_{t}$ are the mean and standard deviation of the time series for a query at time $t$ across repeated downloads. This measures how strong the average signal is compared to its variability. For raw and denoised data, we denote their respective SNR as $\text{SNR}^{\text{(raw)}}_{t}=\hat{\mu}_{t}^{\text{(raw)}}/\hat{\sigma}_{t}^{\text{(raw)}}$ and $\text{SNR}^{\text{(denoised)}}_{t}=\hat{\mu}_{t}^{\text{(denoised)}}/\hat{\sigma}_{t}^{\text{(denoised)}}$. We then compare the two measures by computing the log ratio at each time step $t$, allowing us to track changes in signal quality over time:
$\text{log Ratio of SNR}=\log\left(\frac{\hat{\mu}_{t}^{\text{(denoised)}}}{\hat{\sigma}_{t}^{\text{(denoised)}}}\Big/\frac{\hat{\mu}_{t}^{\text{(raw)}}}{\hat{\sigma}_{t}^{\text{(raw)}}}\right).$ (3)
Figure 9 presents the log ratio of SNR across twenty-seven downloads. Figure 9(a) illustrates its temporal evolution for a single time series. Figure 9(b) displays the mean log ratio for each keyword. Predominantly positive values indicate that denoised data have higher SNR than raw data, confirming that our denoising procedure enhances signal quality by reducing noise and sampling variability.
We further assess the performance of our proposed denoising technique against existing approaches from the literature, in the context of ARIMAX forecasting. Our framework applies smoothing splines iteratively to smooth time series without introducing look-ahead bias. To ensure a fair comparison, we substitute smoothing splines with alternative filters such as moving averages (MA), singular spectrum analysis (SSA), and wavelet transform (WT), while keeping the workflow unchanged. Implementation details are provided in Section S4.1. Figure 10 shows the ratios of MSEs from forecasts using each alternative method relative to those using smoothing splines across all keywords, locations and horizons. The boxplots are mostly above 1, indicating that smoothing splines deliver the most accurate forecasts.
Figure 10: Ratios of MSEs across all locations and horizons for different denoising methods, relative to smoothing splines (benchmark). Values below 1 indicate better performance than smoothing splines.
@@ 5.3 Evaluation of Detrending
Figure 12 summarizes the distribution of $R^{2}$ across all keywords. Fitted trends in the raw data exhibit high $R^{2}$ values (mean =0.77=0.77), confirming that they were well estimated. After detrending, $R^{2}$ values decrease significantly (mean =0.11=0.11), demonstrating the effective removal of deterministic trends. There are a few outliers with low $R^{2}$ values before detrending and high values afterward, suggesting occasional misestimation, as stationarity is detected based on the training set and may not fully capture time series patterns.
@@ 6 Application to Influenza Prediction
For a real-world validation of our preprocessing methodology, we apply the pipeline to the task of real-time influenza forecasting several weeks ahead using both statistical and machine learning models. We retrospectively evaluate state-level forecasts of U.S. influenza hospitalizations (including DC) across two seasons, from October 17, 2022, to April 27, 2024, with training data extending up to October 2022. The application combines Google Trends data of flu-related search terms with hospitalization counts from the U.S. Centers for Disease Control and Prevention (CDC), demonstrating the practical utility of the proposed methodology.
6.1 Identifying Flu-related Terms from Google Trends
In this study, we analyze weekly Google Trends’ “interest over time” data for the United States at the state level since 2004. Our initial set of 161 flu-related terms and topics is based on the work of Yang et al. (2015), identified through Google Correlate in 2009 and 2010, but the service was discontinued in 2019. Therefore, to expand our pool of keywords, we retrieve queries related to “flu”. For long-term seasonal analysis, “top” queries are preferred over “rising” ones, as they exhibit more consistent search patterns in the given time frame. We choose the “Cold and Flu” category to limit results to the area of infectious diseases, and select the entire period from 2004 onward to exclude event-specific queries (e.g., “flu 2024”). Keywords are tailored to each state by setting the corresponding location, which also restricts results to English (or the predominant local language), as opposed to collecting worldwide data. In summary, category, time period, and location serve as filters to refine the dataset. This approach yields time series of weekly search volumes for approximately 500 terms per state, covering symptoms, treatments, medicine, or types of flu.
6.2 Influenza Hospitalizations Data from CDC
The target variable in our case study is influenza-related hospitalizations from the US Centers for Disease Control and Prevention. Weekly admission counts have been collected at the state level in the United States since February 2, 2022 (Mathis et al., 2024) by the National Healthcare Safety Network (NHSN) (Centers for Disease Control and Prevention, 2025c). This data is publicly available through CDC’s GitHub repository (CDC Epidemic Prediction Initiative, 2025). To extend the training period for modeling, we follow the methodology of Meyer et al. (2025) to augment the dataset using ILINet data (Centers for Disease Control and Prevention, 2025b) and FluSurv-Net data (Centers for Disease Control and Prevention, 2025a) with transfer learning. This augmentation provides continuous time series of hospitalization data starting from 2012. Influenza-related hospitalization data is released weekly with a one-week delay, while Google Trends data is available in real time, offering a timely informational advantage when used as exogenous variables in forecasting.
6.3 Modeling
For linear statistical approaches, we consider ARIMAX (Box and Jenkins, 1970; Tiao and Box, 1981), seasonal ARIMAX (Box and Jenkins, 1970; Tiao and Box, 1981), and ARGO (AutoRegression with GOogle search data) (Yang et al., 2015), all of which are widely used and well-established for influenza forecasting (Ray and Reich, 2018; Kandula and Shaman, 2019; Reich et al., 2019). For nonlinear machine learning methods, we apply the tree-based methods LightGBM (Light Gradient Boosting Machine) (Ke et al., 2017) and AdaBoost (adaptive boosting) (Freund and Schapire, 1997), both of which have demonstrated strong performance in influenza prediction (Meyer et al., 2025; Ray et al., 2025; Santillana et al., 2015). Implementation details are provided in Section S4.2
We augment each model with Google Trends data to evaluate the contribution of our preprocessing approach to forecasting performance. Each method is applied separately for every location to produce retrospective out-of-sample nowcasts and 1–3 week-ahead forecasts, simulating real-time predictions. This setup allows us to compare forecast performance across various scenarios: using preprocessed data, topics only, non-preprocessed data, and omitting exogenous variables as baselines.